The second side of a triangle is three times the length of the first. The third side is five less than the second side. If the perimeter is 72, how long is the longest side?
Answer by mananth(16946)

let the first side =x
second = 3x
third side = 3x-5
..
x+3x+3x-5=72
7x-5=72
add 5
7x=72+5
7x=77
x= 11 the first side
second side = 3x= 33
third side = 33-5= 28
...
longest = second side = 33
...
m.ananth@hotmail.ca